The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
                         http://www.astro.uni-bonn.de/~mcnews/

No. 63 February 25, 2002

Contents

Abstracts of 5 refereed papers  1
Abstracts of 3 conference contributions 5

Abstracts of Refereed Papers

Evolution and Geometry of Hot Spots in Supernova Remnant 1987A

Ben E. K. Sugerman^{1}, Stephen S. Lawrence^{2}, Arlin P. S. Crotts^{1}, Patrice Bouchet^{3}, & Steve R. Heathcote^{3}

1 Dept. of Astronomy,Columbia University, New York, NY 10027, USA 2 Dept. of Physics, Hoffstra University, Hempstead, NY 11549, USA 3 CTIO, Casilla 603, La Serena, Chile

We present ground-based near-infrared imaging and HST optical imaging and spectroscopy of the interaction between the ejecta of SN 1987A and its equatorial circumstellar ring. This interaction has made a transition, from emission originating in just a few “hot spots” at restricted locations in position angle around the ring, to a collision producing optical emission over a nearly continuous distribution, with few breaks larger than 45 degrees. The centroids of the first three spots are measured to move at 2000 − 3000 km s^{−1}, which we interpret as a lower limit of the velocity of the forward blast front. Multi-wavelength light curves of the spots show that they do not evolve uniformly, and change significantly on timescales as short as one month; in particular the first spot shows a significant break in its lightcurve. Implications of observed delays between spots appearances are discussed, which leads to a generalized model of hot spot evolution, and suggests that the early appearance of the first hot spot is explained by its inward radial position and a fairly uniform forward blast wave, rather than extraordinary physical circumstances. Data further suggest that the forward blast is reaching the bulk of the inner ring material to the east, the density of which appears higher than elsewhere in the ring. We study the ring geometry, finding evidence suggestive of an intrinsic ellipticity of 0.95, and find lower and upper distance limits of 47.9 ± 0.92 kpc and 54.4 ± 2.1 kpc, respectively.

Accepted by: The Astrophysical Journal

For preprints, contact ben@astro.columbia.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/0202289

The population of massive X-ray binaries. I: The Large Magellanic Cloud

I. Negueruela^{1} and M. J. Coe^{2}

1 Observatoire de Strasbourg, 11 rue de l’Université,F67000 Strasbourg, France 2 Physics and Astronomy Department, Southampton University, Southampton SO17 1BJ, U.K.

We present high resolution blue spectroscopy of an almost complete sample of optical counterparts to massive X-ray binaries in the Large Magellanic Cloud (LMC) and derive their spectral classification. We find an spectral type B0II for the optical counterpart to RX J0532.5−6551, confirming it as the first wind-fed massive X-ray binary in the LMC. We also confirm the Be nature of the proposed counterpart to RX J0535.0−6700. The proposed optical counterpart to RX J0531.5−6518 is a B2V star with signs of emission in the Balmer lines. In total, we give accurate spectral types for 14 counterparts. We find that the overall observed population of massive X-ray binaries in the LMC has a distribution not very different from the observed Galactic population and we discuss different selection effects affecting our knowledge of this population. The spectral distribution of the Be/X-ray binary population is also rather similar to the Galactic one. This distribution implies that Be/X-ray binaries must have preferentially formed from moderately massive binaries undergoing semi-conservative evolution. The observation of several Be/X-ray binaries with large eccentricities implies then the existence of supernova kicks.

Accepted by: Astronomy & Astrophysics

For preprints, contact ignacio@astro.u-strasbg.fr
Also available from the URL http://ArXiv.org/abs/astro-ph/0201451

Discovery of a new Transient X-ray Pulsar in the Small Magellanic Cloud

S. Laycock^{1}, R. H. D. Corbet^{2,3}, D. Perrodin^{2,3}, M. J. Coe^{1}, F. E. Marshall^{2}, C.
Markwardt^{2,4}

1 Dept. of Physics & Astronomy, Southampton University, SO17 1BJ, UK 2 NASA/GSFC, Greenbelt, MD 20771, USA 3 Universities Space Research Association 4 Dept. of Astronomy, University of Maryland, College Park, MD 20742, USA

Systematic Rossi X-Ray Timing Explorer (RXTE) observations of the Small Magellanic Cloud (SMC) have revealed a previously unknown transient X-ray pulsar with a pulse period of 95s. The 95s pulsar, provisionally designated XTE SMC95, was detected in three Proportional Counter Array (PCA) observations during an outburst spanning 4 weeks in March/April 1999. The pulse profile is double peaked reaching a pulse fraction of ≈0.8. The X-ray spectrum is well represented by an absorbed power-law with a photon index of 1.4 and mean unabsorbed flux of ≥ 8.95×10^{−11} ergs cm^{−2} s^{−1}(3 – 25 keV). The source is proposed as a Be/neutron star system on the basis of its pulsations, transient nature and characteristically hard X-ray spectrum. The 2-10 keV X-ray luminosity implied by our observations is ≥ 2×10^{37}ergs s^{−1} which is consistent with that of normal outbursts seen in Galactic systems. This discovery adds to the emerging picture of the SMC as containing an extremely dense population of transient high mass X-ray binaries.

Accepted by: Astronomy & Astrophysics

For preprints, contact sgtl@astro.soton.ac.uk

Also available from the URL astro-ph/0201468

NGC 1866: a workbench for stellar evolution

R. Barmina^{1}, L. Girardi^{1}, and C. Chiosi^{1}

1 Department of Astronomy, University of Padova, Vicolo dell’Osservatorio 2, 35122 Padova, Italy

NGC 1866 is a young, rich star cluster in the Large Magellanic Cloud. Since the cluster is very well populated both in the main sequence and post main sequence stages, thus providing us with a statistically complete sample of objects throughout the various evolutionary phases of intermediate mass stars, it represents a good laboratory for testing stellar evolutionary models. More precisely, NGC 1866 can be used to discriminate among classical stellar models, in which the extension of the convective regions is fixed by the classical Schwarzschild criterion, from models with overshooting, in which an “extra-mixing” is considered to take place beyond the classical limit of the convective zone. Addressing this subject in a recent work, Testa et al.(1999) reached the conclusion that the classical scheme for the treatment of convection represents a good and sufficient approximation for convective interiors. Using their own data, we repeat here the analysis. First we revise the procedure followed by Testa et al.(1999) to correct the data for completeness, second we calculate new stellar models with updated physical input for both evolutionary schemes, finally we present many simulations of the colour-magnitude diagrams and luminosity functions of the cluster using the ratio of the integrated luminosity function of main sequence stars to the number of giants as the normalization factor of the simulations. We also take into account several possible physical agents that could alter the color-magnitude diagram and the luminosity function: they are unresolved binary stars, dispersion in the age, stochastic effects in the initial mass function. Their effect is analyzed separately, with the conclusion that binary stars have the largest impact. The main result of this study is that the convective overshoot hypothesis (together with a suitable percentage of ) unresolved binaries) is really needed to fully match the whole pattern of data. The main drawback of the classical models is that they cannot reproduce the correct ratio of main sequence to post-main sequence stars.

Accepted by: Astronomy & Astrophysics

For preprints, contact barmina@usm.uni-muenchen.de

HD 304373, the second case of 1O/2O Cepheid in the Galaxy

Marco Beltrame^{1} and Ennio Poretti^{2}

1 Università degli Studi, Dipartimento di Fisica, Milano, Italy 2 Osservatorio Astronomico di Brera, Via Bianchi 46, I-23807 Merate, Italy

We report on the discovery of the second case of a galactic Cepheid pulsating in the the first (1O) and second (2O) radial overtones. Our re-analysis of the ASAS Catalogue detected the simultaneous excitation of two modes in the light curve of ASAS 112843 –5925.7≡HD 304373: P_{1}=0.922405 d and P_{2}=0.743241 d (P_{2}/P_{1}=0.8058). In the Galaxy there was only one similar star so far, i.e., CO Aur (P_{1}=1.78303 d, P_{2}=1.42778 d, P_{2}/P_{1}=0.8008).

The ratio between the periods, the short value of the 1O period and the shape of the 1O light curve makes HD 304373 very similar to the 1O/2O Cepheids in the Magellanic Clouds, and in particular to macho*05:30:11.7 -69:52:02 (P_{1} = 0.932517 d and P_{2}=0.751390 d). The implications of a so close similarity between a galactic 1O/2O pulsator and LMC ones are discussed in terms of importance of the metallicity effects.

Submitted to: Astronomy & Astrophysics Letters

For preprints, contact poretti@merate.mi.astro.it

Abstracts of Non-Refereed Papers

Planetary Nebulae in the Magellanic Clouds: Probing Stellar Evolution and Populations

L. Stanghellini

Space Telescope Science Institute, 3700 San Martin Drive, Baltimore (MD) 21218

This review contains: (1) the scientific motivations for studying Planetary Nebulae in the Magellanic Clouds; (2) a review of this field of study, from the origins to the most recent results, focusing on the papers that have been published since the last IAU Symposium on Planetary Nebulae; (3) a review of the Hubble contribution to the field, from the early results to our own Magellanic Cloud Planetary Nebula program.

Invited review, to appear in: IAU Symposium No. 209, Planetary Nebulae: Their Evolution and Role in the Universe (Canberra, Australia November 19 - 23 2001) For preprints, contact lstanghe@stsci.edu

State transitions in LMC X-3

R. Soria^{1}, M. J. Page^{1}, and K. Wu^{1}

1 Mullard Space Science Laboratory, University College London, Holmbury St Mary, Surrey RH5 6NT, UK

We carried out a multiwavelength study of the black-hole candidate LMC X-3 with XMM-Newton. The system showed a transition to a low-hard state, in which the X-ray spectrum was well fitted by a simple power law. It then returned to a high-soft state, characterised by a strong disk-blackbody component. The line-of-sight absorption column density is < ∼ 4 × 10^{20} cm^{−2}, consistent with the foreground Galactic absorption. This rules out wind accretion. We argue that, despite LMC X-3 being a high-mass X-ray binary, Roche-lobe overflow is the main mechanism of mass transfer. ¿From UV/optical observations in the low-hard state, we determine that the companion is a slightly evolved B5 star with a mass M ≈ 4.5 M_{⊙}. This is indeed consistent with the secondary star being close to filling its Roche lobe.

To appear in: Proceedings of the Symposium “New Visions of the X-ray Universe in the XMM-Newton and Chandra Era”, 26-30 November 2001, ESTEC, The Netherlands For preprints, contact rs1@mssl.ucl.ac.uk

Also available from the URL http://xxx.lanl.gov/pdf/astro-ph/0202015

Galactic Cannibalism: the Origin of the Magellanic Stream

S. T. Maddison, D. Kawata and B. K. Gibson

Centre for Astrophysics & Supercomputing, Swinburne University, Australia

We are in a privileged location in the Universe which allows us to observe galactic interactions from close range – the merger of our two nearest dwarf satellite galaxies, the LMC and SMC. It is important to understand the local merger process before we can have confidence in understanding mergers at high redshift. We present high resolution Nbody+SPH simulations of the disruption of the LMC and SMC and the formation of the Magellanic Stream, and discuss the implications for galaxy formation and evolution.

To appear in: “The Evolution of Galaxies II: Basic Building Blocks”, (2002) ed. M. Sauvage et al. (Kluwer)

For preprints, contact smaddison@swin.edu.au
Also available from the URL http://au.arXiv.org/abs/astro-ph/0201401